OpenStudy (anonymous):

help on homework In this assignment, you will be using a recipe of your choosing and using ratios to convert the recipe to make enough for 50 people. This assignment is worth 50 points. Step 1: Choose a recipe that has at least 4 ingredients. Step 2: List all the ingredients and the amount of each ingredient in the recipe. Also list the number of servings it makes. You do not need to list the directions, only the ingredients. [10 points] Step 3: Choose at least four of the ingredients to convert. For each ingredient, set up a proportion that will find the amount needed of that ingredient to make a recipe for 50 servings. Showing all work, find the amount of each of the four ingredients you chose. [10 points per ingredient = 40 points] Step 4: Submit all work and answers to your teacher for grading.

OpenStudy (anonymous):

do you want us to just do this for you? because it sounds pretty straight forward. this also isn't a math question, unless you are asking how to multiply numbers...

OpenStudy (anonymous):

i just font get math

OpenStudy (anonymous):

well for example if your original recipe only serves 4 people, but you need it to serve 50. then divide 50 by 4 to find the ratio of ingredients 50/4 = 12.5 so your ratio is 1:12.5 that means you need 12.5 times the ingredients to serve 50 people if your original recipe only serves 4

OpenStudy (anonymous):

ok so my recipe is sugar cookies and the ingrediants are 3 cups all-purpose flour 3/4 teaspoon baking powder 1/4 teaspoon salt 1 cup unsalted butter, softened 1 cup sugar 1 egg, beaten 1 tablespoon milk

OpenStudy (anonymous):

so i pick 4 of them right ?

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

okay how many cookies does that make?

OpenStudy (anonymous):

about about 3 dozen-2 1/2 inch

OpenStudy (anonymous):

okay so it makes 36 cookies, and you need 50 so 50/36 = 25/18 or 1.388888 so take 4 of your ingredients and multiply their amounts by 25/18 or 1.388888

OpenStudy (anonymous):

any 4 ?

OpenStudy (anonymous):

@robz8 some of my ingredients asks for 1 so do i just divide 50/1?

OpenStudy (anonymous):

or do i multiply them?